OpenStudy (gabylovesyou):

Factor completely x2 + 20x − 36. (x − 2)(x − 18) (x + 2)(x − 18) (x − 2)(x + 18) Prime I think its prime.. correct ?

OpenStudy (skullpatrol):

Multiply out the other answer choices and find out :-)

OpenStudy (skullpatrol):

Do you know FOIL?

OpenStudy (gabylovesyou):

i don't think it can be factored... so if u can't, its prime right ?

OpenStudy (gabylovesyou):

nd yes i do

OpenStudy (skullpatrol):

Yes, just like an integer: If it can not be factored, then it is prime.

OpenStudy (gabylovesyou):

Thank you

OpenStudy (skullpatrol):

Thanks for asking :-)

OpenStudy (zarkon):

it does factor \[x^2 + 20x - 36=(x+2(5-\sqrt{34}))(x+2(5+\sqrt{34}))\] it just doesn't factor over the integers

OpenStudy (gabylovesyou):

ohhh okkk tyty :D

OpenStudy (skullpatrol):

A polynomial cannot be factored over the set of polynomials with the integer coefficients is called a prime polynomial.

OpenStudy (skullpatrol):

An integer that cannot be factored over the set of integers is called a prime number.